OpenStudy (anonymous):

HELP!! 1.) 2x+y=-1 x+ 4y=-10 By substituting. 2.) 2x+y=-1 3x-y=16 By graphing.

OpenStudy (anonymous):

Find the values for x from second equation first..

OpenStudy (anonymous):

1) 2x+y= -1.......(1) x+4y = -10......(2) from (2) x = -10-4y....(3) subst (3) into (1) 2(-10-4y)+y = -1 solve for y

OpenStudy (anonymous):

\(x = -10 - 4y\) Now plug this in first equation : \(2(-10 - 4y) + y = -1\) \(-20 - 8y + y = -1\) \(-7y = 19\) Find y from here and then find x by plugging the found value of y in any one equation..

OpenStudy (anonymous):

graphing you would need a graph paper. sketching the graph is not enough. convert the equations into y = mx +c format m is the gradient and c is the vertical intercept. the intersection points are the solutions for x and y

OpenStudy (anonymous):

@irkiz Is the answer suppose to be a decimal???

OpenStudy (anonymous):

For y you can simply write : \[y = - \frac{19}{7}\]

OpenStudy (anonymous):

@waterineyes i tried that already and it's not right!

OpenStudy (anonymous):

Let me check once again then..

OpenStudy (anonymous):

y is coming right according to me..

OpenStudy (anonymous):

so y = -2.11 ?

OpenStudy (anonymous):

Yes it is..

OpenStudy (anonymous):

NO..

OpenStudy (anonymous):

It is : y = -2.71

OpenStudy (anonymous):

You know what thanks for the help but nevermind because i keep trying to figure it out & it's not happening because i keep getting -2.111111111111. So idk...